# Revision of conservative lower bound on intergalactic magnetic field from Fermi and Cherenkov telescope observations of extreme blazars

### v1

@@ 1 Introduction
Very-High-Energy (VHE) $\gamma$-rays from distant blazars (a type of Active Galactic Nuclei with jets aligned along the line of sight) interact with optical-infrared Extragalactic Background Light (EBL) photons on their way from the source to the Earth and deposit electron-positron pairs in the intergalactic medium, mostly in the voids of the Large Scale Structure. The electrons and positrons subsequently loose energy via inverse Compton scattering of photons of Cosmic Microwave Background and EBL and produce secondary $\gamma$-ray emission detectable by $\gamma$-ray telescopes (Aharonian et al., 1994). In the absence of magnetic fields in the voids, the secondary $\gamma$-ray flux would be emitted in the direction of the primary $\gamma$-ray beam and provide an additional contribution to the blazar $\gamma$-ray flux. Inter-Galactic Magnetic Field (IGMF) present in the voids deflects the trajectories of electrons and positrons so that the secondary emission is not any more directed along the primary $\gamma$-ray beam, suppressing the secondary flux in an IGMF-dependent way (Plaga, 1995; Neronov & Semikoz, 2007). This suppression has been observed (Neronov & Vovk, 2010) and used to establish the existence of non-zero magnetic field in the voids (Neronov & Vovk, 2010; Taylor et al., 2011; Acciari et al., 2023a; Ackermann et al., 2018; Aharonian et al., 2023).
@@
Aharonian et al. (2023) have considered a combination of Fermi/LAT data with the data of HESS IACT for five blazars and derived lower bounds on IGMF, gradually relaxing the assumptions about the activity cycles of those sources: decade long $T\sim 10$ yr (similar to Acciari et al. (2023a)), $T\sim 10^{4}$ yr or $T\sim 10^{7}$ yr. The bounds derived by Aharonian et al. (2023) are much stronger, in particular for the ”conservative” time scale $T\sim 10$ yr, $B>7\times 10^{16}$ G. In the specific case of 1ES 0229+200, the lower bound derived by Aharonian et al. (2023) is $B>4\times 10^{-16}$ G, a factor-of-twenty better than that of Acciari et al. (2023a). This is surprising, given that the analysis of Acciari et al. (2023a) includes also most of the HESS data considered by Aharonian et al. (2023).
The flux of the secondary $\gamma$-ray emission from the intergalactic medium is determined by the level of the primary VHE source flux in multi-TeV energy range. This flux is maximized for the sources with hard VHE $\gamma$-ray spectra and high VHE flux extending without high-energy cut-off in the TeV range. This source class is known as ”extreme” blazars (Fossati et al., 1998). The VHE emission from the extreme blazars is believed to be produced via synchrotron-self-Compton mechanism, via up-scattering of the synchrotron photons by energetic electrons. The frequency of the high-energy cut-off in the synchrotron spectrum correlates with the energy of the cut-off in the $\gamma$-ray spectrum. The ”extreme” blazars are those with the highest energy synchrotron cut-off frequencies (in excess of $10^{17}$ Hz) (Chang et al., 2019). The extreme blazars detected in the VHE band, the ”extreme TeV blazars” (Biteau et al., 2020), are the best targets for the IGMF analysis.
Neronov & Semikoz (2025) have compiled a catalog of VHE emitting AGN based on a long exposure of Fermi/LAT. This catalog lists 275 such AGN in the sky region at Galactic latitudes $|b|>10^{\circ}$. The catalog includes 63 extreme blazars. 22 of them have already been detected by IACTs.
In what follows we consider the IACT detected extreme blazars to re-evaluate the ”conservative” lower bound on IGMF. In parallel, we are able to clarify the origin of discrepancy between the bound derived by Acciari et al. (2023a) and Aharonian et al. (2023). Our analysis is an improvement on both works, because it profits from a better knowledge of the ”best targets” for the IGMF analysis, that were not considered in these two works.
@@ 2 VHE detected extreme blazar set
Table 1 presents the list of extreme blazars detected in the VHE band by both Fermi/LAT and IACTs (Neronov & Semikoz, 2025). As mentioned in the introduction, there are further 41 VHE detected extreme blazars reported by Neronov & Semikoz (2025), for which there is currently no IACT data available. For these sources we cannot yet judge if they provide promising targets for the IGMF analysis and we do not consider them in this work. We also do not consider the brightest source, Mrk 501, for which dedicated studies have found that the secondary flux is largely sub-dominant (Takahashi et al., 2012; Korochkin et al., 2021).
We have combined broad-band GeV-TeV spectra of the 21 sources from Table 1 by complementing Fermi/LAT measurements with measurement by IACTs reported in the literature (cited in the ”Time” column of the table). To extract Fermi/LAT spectra we use the aperture photometry approach in the same way as Neronov & Semikoz (2025) (see Appendix A for details). We fit all the combined Fermi/LAT $+$ IACT spectra with cut-off powerlaw model $N(E/1\mbox{ TeV})^{-\Gamma}\exp(-E/E_{cut})$ modified by the effect of absorption on EBL, taken from Saldana-Lopez et al. (2021). We derive values of the slope $\Gamma$, cut-off energy $E_{cut}$ and normalization of the spectrum $N$ for each source. The values of $\Gamma,E_{cut}$ are given in columns 5 and 6 of Table 1. For most of the extreme blazars, the spectrum extends well into the TeV energy range and we are only able to obtain a lower bound on the cut-off energy, as can be seen from the ”$E_{cut}$” column of the table.
@@
(1) H. E. S. S. Collaboration et al. (2013); (2) Aharonian et al. (2023); (3) Madhavan & VERITAS Collaboration (2013); (4) Acciari et al. (2023b); (5) Aliu et al. (2012); (6) Orr (2013); (7) Acciari et al. (2010); (8) Collaboration (2019); (9) Acciari et al. (2020); (10) Bony et al.(2022)Bony (de), Bylund, Meyer, Noel, Sanchez, Abdalla, Aharonian, Ait-Benkhali, Anguener, Arcaro, Armand, Armstrong, Ashkar, Backes, Baghmanyan, Barbosa Martins, Barnacka, Barnard, Batzofin, Becherini, Berge, Bernloehr, Bi, Böttcher, Boisson, Bolmont, Breuhaus, Brose, Brun, Bulik, Cangemi, Caroff, Casanova, Catalano, Chambery, Chand, Chen, Cotter, Curlo, Dalgleish, Damascene Mbarubucyeye, Davids, Davies, Devin, Djannati-Ataï, Dmytriiev, Donath, Doroshenko, Dreyer, Du Plessis, Duffy, Egberts, Einecke, Ernenwein, Fegan, Feijen, Fiasson, Fichet de Clairfontaine, Fontaine, Frans, Fuessling, Funk, Gabici, Gallant, Giavitto, Giunti, Glawion, Glicenstein, Grondin, Hattingh, Haupt, Hermann, Hinton, Hofmann, Hoischen, Holch, Holler, Horns, Huang, Huber, Hörbe, Jamrozy, Jankowsky, Joshi, Jung, Kasai, Katarzynski, Katz, Khangulyan, Khelifi, Klepser, Kluzniak, Komin, Konno, Kosack, Kostunin, Kreter, Kukec Mezek, Kundu, Lamanna, Le Stum, Lemiere, Lemoine-Goumard, Lenain, Leuschner, Levy, Lohse, Luashvili, Lypova, Mackey, Majumdar, Malyshev, Malyshev, Marandon, Marchegiani, Marcowith, Mares, Marti’i-Devesa, Marx, Maurin, Meintjes, Mitchell, Moderski, Mohrmann, Montanari, Moore, Morris, Moulin, Muller, Murach, Nakashima, Naurois (de), Nayerhoda, Davids, Niemiec, O’Brien, Oberholzer, Ohm, Olivera-Nieto, Ona-Wilhelmi (de), Ostrowski, Panny, Panter, Parsons, Peron, Pita, Poireau, Prokhorov, Prokoph, Puehlhofer, Punch, Quirrenbach, Reichherzer, Reimer, Reimer, Remy, Renaud, Reville, Rieger, Romoli, Rowell, Rudak, Rueda Ricarte, Ruiz Velasco, Sahakian, Sailer, Salzmann, Santangelo, Sasaki, Schaefer, Schutte, Schwanke, Schüssler, Senniappan, Seyffert, Shapopi, Shiningayamwe, Simoni, Sinha, Sol, Spackman, Specovius, Spencer, Spir-Jacob, Stawarz, Steenkamp, Stegmann, Steinmassl, Steppa, Sun, Takahashi, Tanaka, Tavernier, Taylor, Terrier, Thiersen, Thorpe-Morgan, Tluczykont, Tomankova, & Tsirou (de); (11) Hayashida et al. (2016); (12) HESS Collaboration et al. (2013); (13) Ribeiro (2023); (14) Farrell (2022); (15) Archambault et al. (2016); (16) Cao et al. (2024); (17) Abeysekara et al. (2016).
@@ 3 Secondary emission from the intergalactic medium
The cut-off powerlaw fit to the source spectra corresponds to the physical model with ”infinite” IGMF, i.e. with IGMF field strong enough to completely suppress the secondary flux from the direction of the source. Another extreme is the case with zero IGMF. For each source, we attempt to fit such zero IGMF alternative model to the spectrum. In this model, the total flux is a sum of the primary source flux attenuated by the effect of pair production on EBL plus the secondary flux from the inverse Compton scattering by the electron positron pairs. In the absence of IGMF, the secondary flux is emitted in the same direction as the primary source flux and is contributing to the point source spectrum. There is also (almost) no time delay of the secondary signal, so that even if all the sources from Table 1 have only appeared on the sky 20 years ago (with the start of observations by IACTs), their signals should be accompanied by the secondary emission component. In a more general situation, the secondary $\gamma$-rays arrive with a delay (Neronov & Semikoz, 2009)
@@
To move beyond this order-of-magnitude estimate, we use CRbeam code for modeling of the secondary emission (Berezinsky & Kalashev, 2016; Kalashev et al., 2023). We consider IGMF with correlation length $\lambda_{B}=1$ Mpc $>l_{e}$ in the energy range of interest (Neronov & Semikoz, 2009). The magnetic field is chosen to have Kolmogorov power spectrum down to the wave number $k_{min}=2\pi/(5\lambda_{B})$. We assume that the source switches on as a $\theta$-function at the start of the 10-year long interval and calculate the average secondary signal spectrum over the 10-year time period from the start of the source activity. This secondary spectrum is added to the attenuated primary source spectrum and the combined primary $+$ secondary model spectrum is fit to the data.
The time delay $T_{d}$ is related to the angle of misalignment of the arrival direction of the secondary $\gamma$-ray $\theta$ as $T_{d}\sim\theta^{2}D$, where $D$ is the distance to the $\gamma$-ray source (Neronov & Semikoz, 2009). The time delay of the order of $20$ yr corresponds to the misalignment angle $\theta\sim 10^{\prime\prime}$, which is not resolvable with $\gamma$-ray telescopes. Given this fact, our model fitting is limited to the point source spectrum.
We compare the quality of fits of the data with the two alternative models (zero IGMF and infinite IGMF), using the $\chi^{2}$ minimization. Table 1 provides the $\chi^{2}$ values of the fits of the data with the cut-off powerlaw model of the intrinsic spectrum (i.e. with very large IGMF that completely suppresses the secondary flux), $\chi^{2}_{\infty}$ and with the zero IGMF model, $\chi^{2}_{0}$. The $\chi^{2}$ values are quoted together with the number of degrees of freedom (the number of spectral data points minus 3, which is the number of parameters of the cut-off powerlaw model). We judge the source useful for the IGMF analysis if the $B=0$ model is found to be inconsistent with the data. Considering the magnetic field strength as an additional parameter, we consider the model with zero magnetic field inconsistent with the data if $\Delta\chi^{2}=\chi^{2}_{0}-\chi^{2}_{\infty}>2.71$.
@@ 3.1 Sources that do not yield constraints on IGMF
For most of the sources in Table 1 we were able to find an acceptable fit to the Fermi/LAT $+$ IACT data using the $B=0$ model. We show spectra of these sources in Fig. 6 in the Appendix. The spectrum of the primary source is shown by the dashed line and the secondary flux spectrum is shown by the dotted line. The sum of the attenuated primary source and the secondary source fluxes is shown by the solid line. As an example, the spectrum of the first source, SHBL J001355.9-185406, has a high-energy cut-off. This is the reason why the secondary flux expectation is low. The level of the secondary flux is factor-of-two lower than the primary source flux at GeV energy and a small adjustment of the slope of the primary source spectrum (compared to the model without the secondary flux component) is sufficient to make the model with the secondary flux consistent with the data. In most of the other cases the cut-off in the spectrum is not constrained, but the quality of the VHE data is insufficient for a good estimate of the secondary flux level. It is clear that improvement of the quality of the VHE spectral measurements (e.g. via longer exposure with existing IACTs, or with the next-generation IACTs) can tighten the constraint on the intrinsic source spectrum shape, in particular, on the lower bound on $E_{cut}$. If the lower bound on $E_{cut}$ moves toward higher energy, the level of the secondary flux increases and it can start to be too high to be compatible with Fermi/LAT measurements. This would make these sources useful for the IGMF analysis. If the improvement of the IACT data would yield measurement of $E_{cut}$, precise predictions for the secondary flux level and spectra shape would become possible, further improving the precision of the IGMF analysis.
Several sources that have been previously used to derive constraints on IGMF by Aharonian et al. (2023) are not constraining IGMF in our analysis, in spite of the fact that we are using the same Fermi/LAT $+$ HESS data combination as Aharonian et al. (2023). From Fig. 6 one can see that, contrary to Aharonian et al. (2023), we find that the spectra of these sources can be fit consistently with a model with zero IGMF. We explain the origin of this discrepancy in Appendix A.
@@ 3.2 Sources that yield constraints on IGMF
There are seven sources for which the fit with zero IGMF is incompatible with the data: 1ES 0229+200, 1ES 0347-121, 1ES 0502+675, 1ES 1101-232, H 1426+428, 1ES 1727+502 and H 2356-309. Three of the sources: 1ES 0229+200, 1ES 0347-121 and 1ES 1101-232, have been used by Neronov & Vovk (2010) in the initial analysis that resulted in inference of existence of non-zero IGMF in the voids of the Large Scale Structure. H 2356-309 has been considered in the IGMF search by Aharonian et al. (2023). Three other sources, 1ES 0502+675, 1ES 1727+502 and H 1426+428 have not yet been considered in the IGMF analysis.
@@
Fig. 3 shows the $\Delta\chi^{2}=\chi^{2}(B|E_{cut},\Gamma,N)-\chi^{2}_{\infty}$ of the model fits as a function of $B$, with the intrinsic spectrum parameters $E_{cut},\Gamma,N$ fixed to their best-fit values for each $B$. Following Avni (1976), we assume that $\chi^{2}(B)=\chi^{2}(B|E_{cut},\Gamma,N)$ is distributed as $\chi^{2}_{1}$ with one degree of freedom and estimate one-sided 95% lower bound on $B$ by requiring $\Delta\chi^{2}<2.71$ for the models with acceptable $B$. In this sense, our approach is equivalent to that of Aharonian et al. (2023) where the same principle was applied to the profile likelihood as a function of $B$.
@@
stems from the spectral fit to the data of 1ES 0502+675 (Fig. 3). Neronov & Semikoz (2025) have noticed that this is the second-brightest VHE detected extreme blazar after Mrk 501. It has hard $\gamma$-ray spectrum extending up to TeV and as such it is a better candidate for constraining IGMF. Our analysis confirms this expectation. The IACT data considered in our analysis are from Orr (2013). From Fig. 1 one can see that the combined Fermi/LAT+VERITAS source spectrum does not show a signature of high-energy cut-off, once corrected for the EBL. The spectrum is hard, with the slope $\Gamma$ consistent with the hardest possible slope of inverse Compton emission (1.5, see Table 1). The fit with minimal possible IGMF features even harder intrinsic powerlaw slope, $\Gamma=1.3$. Table 2 lists the constraints on $B$ imposed by the data on the other six sources. Fig. 3 shows the $\Delta\chi^{2}$ profiles for all the seven sources.
@@
Remarkably, the source 1ES 0229+200 that was found to yield strongest constraints in the HESS $+$ Fermi/LAT analysis by Aharonian et al. (2023) is found to yield an order-of-magnitude weaker bound compared to Aharonian et al. (2023), in spite of the fact that the same HESS data are used in our analysis. We explain in Appendix A that this difference stems from the mismatches in the modeling of the secondary flux in the two analyses. Our secondary flux model rather matches that of Acciari et al. (2023b). The spectrum of the model with minimal possible IGMF in this case is shown in the left column of the top panel of Fig. 2.
@@
Our bound on IGMF from 1ES 0229+200 data is broadly consistent with that derived by Acciari et al. (2023a) based on a combination of Fermi/LAT data with those of MAGIC, VERITAS and HESS. Acciari et al. (2023a) have noticed that 1ES 0229+200 is variable on the time scale of 20 yr of historical observations with IACTs. This variability can be taken into account self-consistently, by modeling the lightcurve of the delayed secondary flux using the knowledge of the primary source lightcurve. In our simplified version of analysis, we do not consider the details of the primary source variability. This explains the fact that our bound from 1ES 0229+200 data is somewhat weaker than that of Acciari et al. (2023a).
@@
A bound similar to that from the 1ES 0229+200 data is provided by the 1ES 1101-232 data (see Fig. 3 and Table 2). Also this bound is significantly lower than the bound derived by Aharonian et al. (2023), with the discrepancy stemming from the modeling of the secondary signal. Four more sources: 1ES 0347-121, H 1426-428, 1ES 1727+502 and H 2356-309, provide lower bounds on IGMF that are much weaker than that imposed by the 1ES 0502+675 modeling. Weakness of the lower bound from 1ES 0347-121 can be partially understood as stemming from the bad quality of the spectral fits. Table 1 indicates that the $\chi^{2}_{\infty}$ of the fit of the intrinsic source spectrum with a cut-off powerlaw model is 29.1 for 16 degrees-of-freedom (number of spectral measurement points minus the number of the model parameters). This means that the cut-off powerlaw model is actually inconsistent with the data at $>97\%$ confidence level. The spectrum of the source is shown in the top rigtht panel of Fig. 2. One can see that the spectral data show large deviations from the powerlaw behavior and also HESS data are not matching well the highest energy data points of Fermi/LAT. This can be because of variability of the source: Fermi/LAT measures the time-average source flux, while HESS performs snapshot observations on an irregular basis, that can catch some activity episodes and miss quiescent periods. In fact, the fit favors the model with magnetic field somewhat below $10^{-17}$ G, as can be seen from the $\Delta\chi^{2}$ plot for this source in Fig. 3. In this case the ”hump” in the spectrum provided by the secondary flux component fits one of the deviations from the powerlaw in the data. The same is true for H 1426+428. The source spectrum is shown in the right column in the second row of Fig. 2. The fit favors certain value of $B$ for which the secondary flux component fits a hump in the Fermi/LAT spectrum.

### v2

@@ 1 Introduction
Very-High-Energy (VHE) $\gamma$-rays from distant blazars (a type of Active Galactic Nuclei with jets aligned along the line of sight) interact with optical-infrared Extragalactic Background Light (EBL) photons on their way from the source to the Earth and deposit electron-positron pairs in the intergalactic medium, mostly in the voids of the Large Scale Structure. The electrons and positrons subsequently loose energy via inverse Compton scattering of photons of Cosmic Microwave Background and EBL and produce secondary $\gamma$-ray emission detectable by $\gamma$-ray telescopes (Aharonian et al., 1994). In the absence of magnetic fields in the voids, the secondary $\gamma$-ray flux would be emitted in the direction of the primary $\gamma$-ray beam and provide an additional contribution to the blazar $\gamma$-ray flux. Inter-Galactic Magnetic Field (IGMF) present in the voids deflects the trajectories of electrons and positrons so that the secondary emission is not any more directed along the primary $\gamma$-ray beam, suppressing the secondary flux in an IGMF-dependent way (Plaga, 1995; Neronov and Semikoz, 2007). This suppression has been observed (Neronov and Vovk, 2010) and used to establish the existence of non-zero magnetic field in the voids (Neronov and Vovk, 2010; Taylor et al., 2011; Acciari et al., 2023a; Ackermann et al., 2018; Aharonian and others, 2023).
@@
Aharonian and others (2023) have considered a combination of Fermi/LAT data with the data of HESS IACT for five blazars and derived lower bounds on IGMF, gradually relaxing the assumptions about the activity cycles of those sources: decade long $T\sim 10$ yr (similar to Acciari et al. (2023a)), $T\sim 10^{4}$ yr or $T\sim 10^{7}$ yr. The bounds derived by Aharonian and others (2023) are much stronger, in particular for the ”conservative” time scale $T\sim 10$ yr, $B>7\times 10^{16}$ G. In the specific case of 1ES 0229+200, the lower bound derived by Aharonian and others (2023) is $B>4\times 10^{-16}$ G, a factor-of-twenty better than that of Acciari et al. (2023a). This is surprising, given that the analysis of Acciari et al. (2023a) includes also most of the HESS data considered by Aharonian and others (2023).
The flux of the secondary $\gamma$-ray emission from the intergalactic medium is determined by the level of the primary VHE source flux in multi-TeV energy range. This flux is maximized for the sources with hard VHE $\gamma$-ray spectra and high VHE flux extending without high-energy cut-off in the TeV range, provided that the $\gamma\gamma$ opacity of the EBL is sufficient to convert primary photons into secondary pairs. The latter may result in a negligible cascade component for nearby sources despite having a hard primary spectrum. Sources exhibiting these spectral characteristics are known as ”extreme” blazars (Fossati et al., 1998). The VHE emission from the extreme blazars is believed to be produced via synchrotron-self-Compton mechanism, via up-scattering of the synchrotron photons by energetic electrons. The frequency of the high-energy cut-off in the synchrotron spectrum correlates with the energy of the cut-off in the $\gamma$-ray spectrum. The ”extreme” blazars are those with the highest energy synchrotron cut-off frequencies (in excess of $10^{17}$ Hz) (Chang et al., 2019). The extreme blazars detected in the VHE band, the ”extreme TeV blazars” (Biteau et al., 2020), are the best targets for the IGMF analysis.
Neronov and Semikoz (2025) have compiled a catalog of VHE emitting AGN based on a long exposure of Fermi/LAT. This catalog lists 275 such AGN in the sky region at Galactic latitudes $|b|>10^{\circ}$. The catalog includes 63 extreme blazars. 22 of them have already been detected by IACTs.
In what follows we consider the IACT detected extreme blazars to re-evaluate the ”conservative” lower bound on IGMF. In parallel, we are able to clarify the origin of discrepancy between the bound derived by Acciari et al. (2023a) and Aharonian and others (2023). Our analysis is an improvement on both works, because it profits from a better knowledge of the ”best targets” for the IGMF analysis, that were not considered in these two works.
@@ 2 VHE detected extreme blazar set
Table 1 presents the list of extreme blazars detected in the VHE band by both Fermi/LAT and IACTs (Neronov and Semikoz, 2025). As mentioned in the introduction, there are further 41 VHE detected extreme blazars reported by Neronov and Semikoz (2025), for which there is currently no IACT data available. For these sources we cannot yet judge if they provide promising targets for the IGMF analysis and we do not consider them in this work. We also do not consider the brightest source, Mrk 501, for which dedicated studies have found that the secondary flux is largely sub-dominant (Takahashi et al., 2012; Korochkin et al., 2021).
We have combined broad-band GeV-TeV spectra of the 21 sources from Table 1 by complementing Fermi/LAT measurements with measurement by IACTs reported in the literature (cited in the ”Time” column of the table). To extract Fermi/LAT spectra we use the aperture photometry approach in the same way as Neronov and Semikoz (2025) (see Appendix A for details). We fit all the combined Fermi/LAT ++ IACT spectra with cut-off powerlaw model $N(E/1\mbox{ TeV})^{-\Gamma}\exp(-E/E_{cut})$ modified by the effect of absorption on EBL, taken from Saldana-Lopez et al. (2021). We derive values of the slope $\Gamma$, cut-off energy $E_{cut}$ and normalization of the spectrum $N$ for each source. The values of $\Gamma,E_{cut}$ are given in columns 5 and 6 of Table 1. For most of the extreme blazars, the spectrum extends well into the TeV energy range, creating a degeneracy in the fit for the cut-off energy after a certain threshold. Consequently, we consider only the lowest cut-off energy compatible with the fit and define it as a lower bound as can be seen from the $E_{cut}$” column of the table.
@@
(1) H. E. S. S. Collaboration et al. (2013); (2) Aharonian and others (2023); (3) Madhavan and VERITAS Collaboration (2013); (4) Acciari et al. (2023b); (5) Aliu et al. (2012); (6) Orr (2013); (7) Acciari et al. (2010); (8) Collaboration (2019); (9) Acciari et al. (2020); (10) Bony (de) et al. (2022); (11) Hayashida et al. (2016); (12) HESS Collaboration et al. (2013); (13) Ribeiro (2023); (14) Farrell (2022); (15) Archambault et al. (2016); (16) Cao et al. (2024); (17) Abeysekara et al. (2016).
@@ 3 Secondary emission from the intergalactic medium
The cut-off powerlaw fit to the source spectra corresponds to the physical model with ”infinite” IGMF, i.e. with IGMF field strong enough to completely suppress the secondary flux from the direction of the source. Another extreme is the case with zero IGMF. For each source, we attempt to fit such zero IGMF alternative model to the spectrum. In this model, the total flux is a sum of the primary source flux attenuated by the effect of pair production on EBL plus the secondary flux from the inverse Compton scattering by the electron positron pairs. In the absence of IGMF, the secondary flux is emitted in the same direction as the primary source flux and is contributing to the point source spectrum. There is also (almost) no time delay of the secondary signal, so that even if all the sources from Table 1 have only appeared on the sky 20 years ago (with the start of observations by IACTs), their signals should be accompanied by the secondary emission component. In a more general situation, the secondary $\gamma$-rays arrive with a delay (Neronov and Semikoz, 2009)
@@
To move beyond this order-of-magnitude estimate, we use CRbeam code for modeling of the secondary emission (Berezinsky and Kalashev, 2016; Kalashev et al., 2023). We consider IGMF with correlation length $\lambda_{B}=1$ Mpc $l_{e}$ in the energy range of interest (Neronov and Semikoz, 2009). The magnetic field is chosen to have Kolmogorov power spectrum down to the wave number $k_{min}=2\pi/(5\lambda_{B})$. We assume that the source switches on as a $\theta$-function at the start of the 10-year long interval and calculate the average secondary signal spectrum over the 10-year time period from the start of the source activity. This secondary spectrum is added to the attenuated primary source spectrum and the combined primary ++ secondary model spectrum is fit to the data.
The time delay $T_{d}$ is related to the angle of misalignment of the arrival direction of the secondary $\gamma$-ray $\theta$ as $T_{d}\sim\theta^{2}D$, where $D$ is the distance to the $\gamma$-ray source (Neronov and Semikoz, 2009). The time delay of the order of $20$ yr corresponds to the misalignment angle $\theta\sim 10^{\prime\prime}$, which is not resolvable with $\gamma$-ray telescopes. Given this fact, our model fitting is limited to the point source spectrum.
We compare the quality of fits of the data with the two alternative models (zero IGMF and infinite IGMF), using the $\chi^{2}$ minimization. Table 1 provides the $\chi^{2}$ values of the fits of the data with the cut-off powerlaw model of the intrinsic spectrum (i.e. with very large IGMF that completely suppresses the secondary flux), $\chi^{2}_{\infty}$ and with the zero IGMF model, $\chi^{2}_{0}$. The $\chi^{2}$ values are quoted together with the number of degrees of freedom (the number of spectral data points minus 3, which is the number of parameters of the cut-off powerlaw model). We judge the source useful for the IGMF analysis if the $B=0$ model is found to be inconsistent with the data. Considering the magnetic field strength as an additional parameter, we consider the model with zero magnetic field inconsistent with the data if $\Delta\chi^{2}=\chi^{2}_{0}-\chi^{2}_{\infty}>2.71$ corresponding to the one-sided 95% confidence level (CL).
@@ 3.1 Sources that do not yield constraints on IGMF
For most of the sources in Table 1 we were able to find an acceptable fit to the Fermi/LAT ++ IACT data using the $B=0$ model. We show spectra of these sources in Fig. 6 in the Appendix. The spectrum of the primary source is shown by the dashed line and the secondary flux spectrum is shown by the dotted line. The sum of the attenuated primary source and the secondary source fluxes is shown by the solid line. As an example, the spectrum of the first source, SHBL J001355.9-185406, has a high-energy cut-off. This is the reason why the secondary flux expectation is low. The level of the secondary flux is factor-of-two lower than the primary source flux at GeV energy and a small adjustment of the slope of the primary source spectrum (compared to the model without the secondary flux component) is sufficient to make the model with the secondary flux consistent with the data. In most of the other cases the cut-off in the spectrum is not constrained, but the quality of the VHE data is insufficient for a good estimate of the secondary flux level. It is clear that improvement of the quality of the VHE spectral measurements (e.g. via longer exposure with existing IACTs, or with the next-generation IACTs) can tighten the constraint on the intrinsic source spectrum shape, in particular, on the lower bound on $E_{cut}$. If the lower bound on $E_{cut}$ moves toward higher energy, the level of the secondary flux increases and it can start to be too high to be compatible with Fermi/LAT measurements. This would make these sources useful for the IGMF analysis. If the improvement of the IACT data would yield measurement of $E_{cut}$, precise predictions for the secondary flux level and spectra shape would become possible, further improving the precision of the IGMF analysis.
One source that has been previously used to derive constraints on IGMF by Aharonian and others (2023), namely PKS 0548-322, is not constraining IGMF in our analysis, in spite of the fact that we are using the same Fermi/LAT ++ HESS data combination as Aharonian and others (2023). From Fig. 6 one can see that, contrary to Aharonian and others (2023), we find that the spectrum of this source can be fit consistently with a model with zero IGMF. We explain the origin of this discrepancy in Appendix A.
@@ 3.2 Sources that yield constraints on IGMF
There are seven sources for which the fit with zero IGMF is incompatible with the data: 1ES 0229+200, 1ES 0347-121, 1ES 0502+675, 1ES 1101-232, H 1426+428, 1ES 1727+502 and H 2356-309. Three of the sources: 1ES 0229+200, 1ES 0347-121 and 1ES 1101-232, have been used by Neronov and Vovk (2010) in the initial analysis that resulted in inference of existence of non-zero IGMF in the voids of the Large Scale Structure. H 2356-309 has been considered in the IGMF search by Aharonian and others (2023). Three other sources, 1ES 0502+675, 1ES 1727+502 and H 1426+428 have not yet been considered in the IGMF analysis.
@@
Fig. 3 shows the $\Delta\chi^{2}=\chi^{2}(B|E_{cut},\Gamma,N)-\chi^{2}_{\infty}$ of the model fits as a function of $B$, with the intrinsic spectrum parameters $E_{cut},\Gamma,N$ fixed to their best-fit values for each $B$. Following Avni (1976), we assume that $\chi^{2}(B)=\chi^{2}(B|E_{cut},\Gamma,N)$ is distributed as $\chi^{2}_{1}$ with one degree of freedom and estimate one-sided 95% lower bound on $B$ by requiring $\Delta\chi^{2}<2.71$ for the models with acceptable $B$. In this sense, our approach is equivalent to that of Aharonian and others (2023) where the same principle was applied to the profile likelihood as a function of $B$.
@@
stems from the spectral fit to the data of 1ES 0502+675 (Fig. 3). Neronov and Semikoz (2025) have noticed that this is the second-brightest VHE detected extreme blazar after Mrk 501. It has hard $\gamma$-ray spectrum extending up to TeV and as such it is a better candidate for constraining IGMF. Our analysis confirms this expectation. The IACT data considered in our analysis are from Orr (2013). From Fig. 1 one can see that the combined Fermi/LAT+VERITAS source spectrum does not show a signature of high-energy cut-off, once corrected for the EBL. The spectrum is hard, with the slope $\Gamma$ consistent with the hardest possible slope of inverse Compton emission (1.5, see Table 1). The fit with minimal possible IGMF features even harder intrinsic powerlaw slope, $\Gamma=1.3$. Table 2 lists the constraints on $B$ imposed by the data on the other six sources. Fig. 3 shows the $\Delta\chi^{2}$ profiles for all the seven sources.
@@
Remarkably, the source 1ES 0229+200 that was found to yield strongest constraints in the HESS ++ Fermi/LAT analysis by Aharonian and others (2023) is found to yield an order-of-magnitude weaker bound compared to Aharonian and others (2023), in spite of the fact that the same HESS data are used in our analysis. We explain in Appendix A that this difference stems from the mismatches in the modeling of the secondary flux in the two analyses. Our secondary flux model rather matches that of Acciari et al. (2023b). The spectrum of the model with minimal possible IGMF in this case is shown in the left column of the top panel of Fig. 2.
@@
Our bound on IGMF from 1ES 0229+200 data is broadly consistent with that derived by Acciari et al. (2023a) based on a combination of Fermi/LAT data with those of MAGIC, VERITAS and HESS. Acciari et al. (2023a) have noticed that 1ES 0229+200 is variable on the time scale of 20 yr of historical observations with IACTs. This variability can be taken into account self-consistently, by modeling the lightcurve of the delayed secondary flux using the knowledge of the primary source lightcurve. In our simplified version of analysis, we do not consider the details of the primary source variability and applied a different assumed time activity of 16 years. Moreover, the best-fit energy cut-off for the IGMF bound inferred from 1ES 0229+200 is $E_{cut}\sim 2.5$ TeV (see Table LABEL:tab:best_fit_params_B_field in appendix), somewhat lower than the one used in Acciari et al. (2023a) where $E_{cut}=7$ TeV, which can lead to a slightly larger suppression of the secondary cascade. This explains the fact that our bound from 1ES 0229+200 data is somewhat weaker than that of Acciari et al. (2023a).
@@
A bound similar to that from the 1ES 0229+200 data is provided by the 1ES 1101-232 data (see Fig. 3 and Table 2). Also this bound is significantly lower than the bound derived by Aharonian and others (2023), with the discrepancy stemming from the modeling of the secondary signal. Four more sources: 1ES 0347-121, H 1426-428, 1ES 1727+502 and H 2356-309, provide lower bounds on IGMF that are much weaker than that imposed by the 1ES 0502+675 modeling. Weakness of the lower bound from 1ES 0347-121 can be partially understood as stemming from the bad quality of the spectral fits. Table 1 indicates that the $\chi^{2}_{\infty}$ of the fit of the intrinsic source spectrum with a cut-off powerlaw model is 29.1 for 16 degrees-of-freedom (number of spectral measurement points minus the number of the model parameters). This means that the cut-off powerlaw model is actually inconsistent with the data at $97\%$ confidence level. The spectrum of the source is shown in the top rigtht panel of Fig. 2. One can see that the spectral data show large deviations from the powerlaw behavior and also HESS data are not matching well the highest energy data points of Fermi/LAT. This can be because of variability of the source: Fermi/LAT measures the time-average source flux, while HESS performs snapshot observations on an irregular basis, that can catch some activity episodes and miss quiescent periods. In fact, the fit favors the model with magnetic field somewhat below $10^{-17}$ G, as can be seen from the $\Delta\chi^{2}$ plot for this source in Fig. 3. In this case the ”hump” in the spectrum provided by the secondary flux component fits one of the deviations from the powerlaw in the data. The same is true for H 1426+428. The source spectrum is shown in the right column in the second row of Fig. 2. The fit favors certain value of $B$ for which the secondary flux component fits a hump in the Fermi/LAT spectrum.